# PTT: Point-Trajectory Transformer for Efficient Temporal 3D Object Detection

### v1

@@ 1 Introduction
Detecting 3D objects [7, 13, 21, 8, 6] is a crucial problem for autonomous driving, enabling vehicles to perceive and interpret their surroundings accurately. With the advent of LiDAR sensors, significant strides have been made in enhancing the precision of 3D object detection. Nevertheless, the inherent sparsity of point cloud data captured by LiDAR sensors poses a significant challenge. The sparse data complicates accurately identifying and localizing objects, particularly in challenging occlusion scenarios.
@@
Nevertheless, these proposal-based techniques suffer from two main limitations. First, sampling object point clouds for each frame in sequential modeling requires a significant memory burden for a long sequence during the training and testing phases. Second, these methods focus on the design of networks for point cloud sequences, which often neglects the significance of the interaction between point and proposal features. In this paper, we propose PTT, a Point-Trajectory Transformer for efficient temporal 3D object detection to handle these two challenges (see Figure 1(b)).
@@
Furthermore, unlike previous approaches that concatenate proposal trajectories with point cloud features, we introduce the point-trajectory transformer with long short-term memory to model the relationship between single-frame point clouds and multi-frame proposals. In this approach, we partition information obtained from multiple frames into long-term and short-term encoding to enhance representations over temporal information. In addition, we infer the future possible trajectory of the proposals to generate future-aware point features. Finally, we design an aggregator to facilitate interactions between point and trajectory features.
@@ 3.2 Analysis of Memory Bank’s Size
Regarding the point cloud of objects, MPPNet [1] with 16 frames (default setting) needs to store per-frame object points, so the memory complexity is $\mathcal{O}(16kN)$, which will increase the complexity based on the frame number. For MSF [5] with 8-frames input, the memory complexity for dealing with the point clouds is $\mathcal{O}(8F)$ since they need to keep the entire sequential point cloud for propagating the proposals into all previous frames to find previous bounding box candidates. Instead, our approach only samples the current-frame object point cloud, which only requires $\mathcal{O}(kN)$ space complexity without being proportional to the frame number, allowing us to leverage longer frames (e.g., 64).
@@ 3.3 Point-Trajectory Feature Generation
Unlike the prior work [1, 5] keeping per-frame objects or entire point clouds, we advocate using single-frame points and multi-frame proposals, a more storage-efficient solution as discussed in Section 3.2. In this section, we propose to encode two kinds of features: Point-to-Proposal and Proposal-to-Proposal features.
Point-to-Proposal Features. Consider the current frame’s points-of-interest ${\mathbf{I}}^{T}$ and the previous $T$-frame box trajectories $\{{\mathbf{B}^{1}},...,{\mathbf{B}^{T}}\}$ for a specific object, our objective is to encode the point features based on historical proposals. We first generate point-to-proposal features based on the offset between points and points of the bounding box (i.e., center and eight corners) $\{{{\mathbf{b}}^{t}_{i}\in{\mathcal{C}}_{3}({\mathbf{B}^{t}})}:i=0,...,8\}$ [18]:
@@
Method Frame Space Complexity
Num RPN Point Proposal
MPPNet [1] 16 $\mathcal{O}(4F)$ $\mathcal{O}(16kN)$ $\mathcal{O}(16kO)$
@@ 3.4 Point-Trajectory Transformer
Figure 3: Point-Trajectory Aggregator. Current-frame point cloud features are squeezed and interact with long-term, short-term, and future memory.
Future-Aware Activation. Future information is also important for 3D estimation in the current frame by improving feature learning. Considering a 3D proposal box with $(x,y,z,w,l,h,\theta)$ with the estimated velocity $(v_{x},v_{y})$ at frame $T$, we can infer the box’s state at future time difference $\Delta t$ as $\mathbf{B}^{T+\Delta t}=(x+\Delta t\cdot v_{x},y+\Delta t\cdot v_{y},z,w,l,h,\theta+\Delta t\cdot v_{\theta})$, where $v_{\theta}$ can be calculated by the difference within previous frames: $v_{\theta}=(\theta_{T}-\theta_{1})/T$.
@@ 4.3 Ablation Study
80.00/78.58 74.31/72.95
81.12/79.78 75.63/74.34
81.15/79.79 75.58/74.27
81.38/80.04 75.88/74.59
Table 7: Ablation study on different attention mechanisms in the proposed point-trajectory aggregator. “SA” and “CA” denote the self-attention and cross-attention operations, respectively.
Effectiveness of the attention mechanisms in the proposed point-trajectory aggregator. Table 7 shows the importance of different attention designs in the point-trajectory aggregator that involves self-attention and cross-attention mechanisms. The results show that simultaneous usage of both attention operations in the proposed point-trajectory aggregator achieves the best performance across all L1 (mAP), L1 (mAPH), L2 (mAP), and L2 (mAPH) metrics.
@@
Frame L1 L2
Length mAP / mAPH mAP / mAPH

### v2

@@ 1 Introduction
Detecting 3D objects [8, 14, 22, 9, 6, 7] is a crucial problem for autonomous driving, enabling vehicles to perceive and interpret their surroundings accurately. With the advent of LiDAR sensors, significant strides have been made in enhancing the precision of 3D object detection. Nevertheless, the inherent sparsity of point cloud data captured by LiDAR sensors poses a significant challenge. The sparse data complicates accurately identifying and localizing objects, particularly in challenging occlusion scenarios.
@@
Nevertheless, these proposal-based techniques suffer from two main limitations. First, sampling object point clouds for each frame in sequential modeling causes a memory burden for a long sequence during the training and testing phases. Second, these methods focus on the design of networks for point cloud sequences, which often neglects the significance of the interaction between point and proposal features. In this paper, we propose PTT, a Point-Trajectory Transformer for efficient temporal 3D object detection to handle these two challenges (see Figure 1(b)).
@@
Furthermore, unlike previous methods that concatenate proposal trajectories with point cloud features, we introduce the point-trajectory transformer with long short-term memory to model the relationship between single-frame point clouds and multi-frame proposals. In this approach, we partition information obtained from multiple frames into long-term and short-term encoding to enhance representations over temporal information. In addition, we infer the future possible trajectory of the proposals to generate future-aware point features. Finally, we design an aggregator to facilitate interactions between point and trajectory features.
@@ 3.2 Analysis of Memory Bank’s Size
Regarding the point cloud of objects, MPPNet [1] with 16 frames (default setting) needs to store each object’s per-frame proxy points $N^{\prime}=N/2$, so the memory complexity is $\mathcal{O}(16kN^{\prime})={O}(8kN)$, which will increase the complexity based on the frame number. For MSF [5] with 8-frames input, the memory complexity for dealing with the point clouds is $\mathcal{O}(8F)$ since they need to keep the entire sequential point cloud for propagating the proposals into all previous frames to find previous bounding box candidates. Instead, our approach only samples the current-frame object point cloud, which only requires $\mathcal{O}(kN)$ space complexity without being proportional to the frame number, allowing us to leverage longer frames (e.g., 64).
@@ 3.3 Point-Trajectory Feature Generation
Unlike the prior work [1, 5] keeping per-frame objects or entire point clouds, we advocate using single-frame points and multi-frame proposals, a more storage-efficient solution as discussed in Section 3.2. We follow MPPNet [1] to infer the previous location of the bounding box based on the estimated velocity for all proposals to link cross-frame boxes. After that, we propose to encode two kinds of features: Point-to-Proposal and Proposal-to-Proposal features.
Point-to-Proposal Features. Consider the current frame’s points-of-interest ${\mathbf{I}}^{T}$ and the latest $T$-frame box trajectories $\{{\mathbf{B}^{1}},...,{\mathbf{B}^{T}}\}$ for a specific object, our objective is to encode the point features based on historical proposals. We first generate point-to-proposal features based on the offset between points and points of the bounding box (i.e., center and eight corners) $\{{{\mathbf{b}}^{t}_{i}\in{\mathcal{C}}_{3}({\mathbf{B}^{t}})}:i=0,...,8\}$ [19]:
@@
Method Frame Num Space Complexity
RPN Point Proposal
MPPNet [1] 16 $\mathcal{O}(4F)$ $\mathcal{O}(8kN)$ $\mathcal{O}(16kO)$
@@ 3.4 Point-Trajectory Transformer
Figure 3: Point-Trajectory Aggregator. Current-frame point cloud features $\mathbf{G}^{T}$ are squeezed and interact with long-term memory $\hat{{\mathbf{M}}}_{l}$, short-term memory $\hat{{\mathbf{M}}}_{s}$, and future memory $\hat{{\mathbf{M}}}^{\prime}_{f}$. See Section 3.4 for more details.
Future-Aware Activation. Inspired by the anticipation of future action in the action recognition field  [29], we observe that future information is also important for 3D estimation in the current frame by improving feature learning. Considering a 3D proposal box with $(x,y,z,w,l,h,\theta)$ with the estimated velocity $(v_{x},v_{y})$ at frame $T$, we can infer the box’s state at future time difference $\Delta t$ as $\mathbf{B}^{T+\Delta t}=(x+\Delta t\cdot v_{x},y+\Delta t\cdot v_{y},z,w,l,h,\theta+\Delta t\cdot v_{\theta})$, where $v_{\theta}$ can be calculated by the difference within previous frames: $v_{\theta}=(\theta_{T}-\theta_{1})/T$.
@@ 4.3 Ablation Study
      ✗       ✗       80.00/78.58       74.31/72.95
      ✓       ✗       81.12/79.78       75.63/74.34
      ✗       ✓       81.15/79.79       75.58/74.27
      ✓       ✓       81.38/80.04       75.88/74.59
Table 7: Ablation study on different attention mechanisms in the proposed point-trajectory aggregator. “SA” and “CA” denote the self-attention and cross-attention operations as shown in Figure 3, respectively.
Effectiveness of the attention mechanisms in the proposed point-trajectory aggregator. Table 7 shows the importance of different attention designs in the point-trajectory aggregator that involves self-attention and cross-attention mechanisms. The results show that simultaneous usage of both attention operations in the proposed point-trajectory aggregator achieves the best performance across all metrics.
@@
      Frame Length       L1       L2
      mAP / mAPH       mAP / mAPH
@@ 5 Conclusions
Acknowledgements.
This work was supported in part by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) contract number 140D0423C0074. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.